In the mythical country of Ecotopia there is no government spending, taxes, or foreign trade. The mpc for the people is .8, c (autonomous consumption) is 50b. and II (Intended Investment) is also 50b. (just accidentally the same).   In a minute, you will fill in the table below, using this data.  (b. stands for billions of dollars) (Recall that AD (me) and AE (textbook) are the same thing, as are II (me) and Ip (textbook). Also recall that C = c + (mpc) x Y and that AD = C + II, so AD = c + ((mpc) x Y) + II. I am sorry, every time I rescale, the bar that is supposed to go over the little c moves and I can’t erase it. So please just remember that little c is actually c-bar and stands for autonomous consumption.  


1
Expert's answer
2021-10-31T18:40:05-0400

MPC=0.8


Investment(II)=50


Autonomous consumption(c)=50 billion.


Marginal propensity to consume is the term used to refer to an additional dollar spent for every additional dollar of income.


Autonomous consumption is the amount of consumption that will always occur,regardless of income.


C=(c)+MPC"\times" GDP


=50+0.8"Y"